# Approximating Discrimination Within Models When Faced With Several Non-Binary Sensitive Attributes

### v1

@@ Abstract
Discrimination mitigation with machine learning (ML) models could be complicated because multiple factors may interweave with each other including hierarchically and historically. Yet few existing fairness measures are able to capture the discrimination level within ML models in the face of multiple sensitive attributes. To bridge this gap, we propose a fairness measure based on distances between sets from a manifold perspective, named as ‘harmonic fairness measure via manifolds (HFM)’ with two optional versions, which can deal with a fine-grained discrimination evaluation for several sensitive attributes of multiple values. To accelerate the computation of distances of sets, we further propose two approximation algorithms named ‘Approximation of distance between sets for one sensitive attribute with multiple values (ApproxDist)’ and ‘Approximation of extended distance between sets for several sensitive attributes with multiple values (ExtendDist)’ to respectively resolve bias evaluation of one single sensitive attribute with multiple values and that of several sensitive attributes with multiple values. Moreover, we provide an algorithmic effectiveness analysis for ApproxDist under certain assumptions to explain how well it could work. The empirical results demonstrate that our proposed fairness measure HFM is valid and approximation algorithms (i.e., ApproxDist and ExtendDist) are effective and efficient.
@@ 1 Introduction
As techniques of machine learning (ML) and deep learning (DL) are flourishing developed and ML/DL systems are widely deployed in real life nowadays, the concern about the underlying discrimination hidden in these models has grown, particularly in high-stakes domains such as healthcare, recruitment, and jurisdiction [1], where equity for all stakeholders is pivotal to prevent unjust outcomes, akin to a discriminatory Matthew Effect. It is of significance to prevent ML models from perpetuating or exacerbating inappropriate human prejudices for not only model performance but also societal welfare. Effectively addressing and eliminating discrimination usually requires a comprehensive grasp of its occurrence, causes, and mechanisms. For instance, a case involving a person changing their gender for lower car insurance rates highlights the complexity of fairness in ML.
@@
Although the impressive practical advancements of ML and DL thrive on abundant data, their trustworthiness and equity heavily hinge on data quality. In fact, one of the primary sources of unfairness identified in the existing literature is biases from the data, possibly collected from various sources such as device measurements and historically biassed human decisions [2]. Moreover, the challenge of data imbalance often looms in human-sensitive domains, amplifying concerns of discrimination and bias propagation in ML models. Then misformed model training would amplify imbalances and biases in data, with wide-reaching societal implications. For example, optimising aggregated prediction errors can advantage privileged groups over marginalised ones. In addition, missing data like instances or values may introduce disparities between the dataset and the target population, leading to biassed results as well. Therefore, in order to ensure fairness and mitigate biases, it is crucial to correctly cope with data imbalance and prevent ML models from perpetuating or even exacerbating inappropriate human prejudices.
@@
To mitigate bias within ML models, the very first step to promptly recognise its occurrence. However, promptly detecting discrimination fully, truly, and faithfully is not quite easy because of plenty of factors interweaving with each other. First, learning algorithms might yield unfair outcomes even with purely clean data due to proxy attributes for sensitive features or tendentious algorithmic objectives. For instance, the educational background of one person might be a proxy attribute for those born in families with a preference for boys. Second, the existence of multiple sensitive attributes and their twist with each other highlight the complexity of bias tackling, like one member from a marginalised group could become one of the majority concerning another factor, or vice versa. Third, dynamic changes and historical factors may need to be taken into account, as bias hidden in data, data imbalance, and present decisions may interweave, causing some interrelated impact and vicious circles. Despite many fairness measures that have been proposed to facilitate bias mitigation, most of them mainly focus on one single sensitive attribute or ones with binary values, and few could handle bias appropriately when facing multiple sensitive attributes with even multiple values. Therefore, it motivates us to investigate a proper tool to deal with bias in such aforementioned scenarios.
@@
In this paper, we investigate the possibility of assessing the discrimination level of ML models in the existence of several sensitive attributes with multiple values. To this end, we introduce a novel fairness measure from a manifold perspective, named ‘harmonic fairness measure via manifolds (HFM)’, with two optional versions (that is, maximum HFM and average HFM). However, the direct calculation of HFM lies on a core distance between two sets, which might be pretty costly. Therefore, we further propose two approximation algorithms that quickly estimate the distance between sets, named as ‘Approximation of distance between sets for one sensitive attribute with multiple values (ApproxDist)’ and ‘Approximation of extended distance between sets for several sensitive attributes with multiple values (ExtendDist)’ respectively, in order to speed up the calculation and enlarge its practical applicable values. Furthermore, we also investigate their algorithmic properties under certain reasonable assumptions, in other words, how effective they could be in achieving the approximation goal. Our contribution in this work is four-fold:
@@
  • We propose a fairness measure named HFM that could reflect the discrimination level of classifiers even simultaneously facing several sensitive attributes with multiple values. Note that HFM has two optional versions, of which both are built on a concept of distances between sets from the manifold perspective.
@@
  • We propose two approximation algorithms (that is, ApproxDist and ExtendDist) that accelerate the estimation of distances between sets, to mitigate its disadvantage of costly direct calculation of HFM.
@@
  • We further investigate the algorithmic effectiveness of ApproxDist and ExtendDist under certain assumptions and provide detailed explanations.
@@ 2.1 Techniques to enhance fairness
Existing mechanisms to mitigate biases and enhance fairness in ML models could be typically divided into three types: pre-processing, in-processing, and post-processing mechanisms, based on when manipulations are applied during model training pipelines. Particularly, recent work on in-processing fairness for DL models mainly falls under two types of approaches: constraint-based and adversarial learning methods [3]. Constraint-based methods usually incorporate fairness metrics directly into the model optimisation objectives as constraints or regularisation terms. For instance, Zemel et al. [4], the pioneer in this direction, put demographic parity constraints on model predictions. Subsequent work also includes using approximations [5] or modified training schemes [6] to improve scalability. Adversarial methods intend to learn representations as fairly as possible by removing sensitive attribute information. In such procedures, additional prediction heads may be introduced for attribute subgroup predictions and the information concerning sensitive attributes would be removed through inverse gradient updating [7, 8] or disentangling features [9, 10, 11, 12]. Other fairness enhancing techniques include data augmentations [13], sampling [14, 15], data noising [16], dataset balancing with generative methods [17, 18, 19], and reweighting mechanisms [20, 21]. Recently, mixup operations [22, 23, 3] are adopted to enhance fairness by blending inputs across subgroups [24, 25]. However, most of these studies focus on protecting one single sensitive attribute and are hardly able to deal with several sensitive attributes all at once. And multi-attribute fairness protection remains relatively rarely explored.
@@ 2.2 Existing fairness metrics and multi-attribute fairness protection
The well-known fairness metrics are generally divided into group fairness—such as demographic parity (DP), equality of opportunity (EO), and predictive quality parity (PQP)—and individual fairness [26, 27, 28, 29, 30]. The former mainly focuses on statistical/demographic equality among groups defined by sensitive attributes, while the latter cares more about the principle that ‘similar individuals should be evaluated or treated similarly.’ However, satisfying fairness metrics all at once is hard to achieve because they are usually not compatible with each other [31]. In practice, it may need to deliberate on the choice of the specified distance in individual fairness [29, 26]. Moreover, the three commonly used group fairness measures (that is, DP, EO, and PQP) can only deal with one single sensitive attribute with binary values. Although extending them to scenarios of one sensitive attribute with multiple values is possible, they are still limited when facing several sensitive attributes at the same time. Recent work includes a newly proposed fairness measure named discriminative risk (DR) [32] that is capable of capturing bias from both individual and group fairness aspects and two fairness frameworks (that is, InfoFair [33] and MultiFair [3]) to deliver fair predictions in face of multiple sensitive attributes. Yet these two fairness frameworks are not measures that could directly evaluate the discrimination level of ML models.
@@ 3 Methodology
In this section, we formally study the measurement of fairness from a manifold perspective. Here is a list of some standard notations we use. In this paper, we denote the scalars by italic lowercase letters (e.g., $x$), the vectors by bold lowercase letters (e.g., $\bm{x}$), the matrices/sets by italic uppercase letters (e.g., $X$), the random variables by serif uppercase letters (e.g., $\mathsf{X}$), the real numbers (resp. the integers, and the positive integers) by $\mathbb{R}$ (resp. $\mathbb{Z}$, and $\mathbb{Z}_{+}$), the probability measure (resp. the expectation, and the variance of one random variable) by $\mathbb{P}(\cdot)$ (resp. $\mathbb{E}(\cdot)$ and $\mathbb{V}(\cdot)$), the indicator function by $\mathbb{I}(\cdot)$, and the hypothesis space (resp. models in it) by $\mathcal{F}$ (resp. $f(\cdot)$).
In this paper, $i\in[n]$ is used to represent $i\in\{1,2,...,n\}$ for brevity. We use $S=\{(\bm{x}_{i},y_{i})\}_{i=1}^{n}$ to denote a dataset where the instances are iid. (independent and identically distributed), drawn from an feature-label space $\mathcal{X\!\times\!Y}$ based on an unknown distribution. The feature/input space $\mathcal{X}$ is arbitrary, and the label/output space $\mathcal{Y}\!=\!\{1,2,...,n_{c}\}(n_{c}\geqslant 2)$ is finite, which could be binary or multi-class classification, depending on the number of labels (i.e., the value of $n_{c}$). Presuming that the considered dataset $S$ is composed of the instances including sensitive attributes, the features of one instance including sensitive attributes $\bm{a}\!=\![a_{1},a_{2},...,a_{n_{a}}]^{\mathsf{T}}$ is represented as $\bm{x}\triangleq(\mathop{\breve{\bm{x}}},\mathop{\bm{a}})$, where $n_{a}\geqslant 1$ is the number of sensitive attributes allowing multiple attributes and $a_{i}\!\in\!\mathbb{Z}_{+}(1\!\leqslant\!i\!\leqslant\!n_{a})$ allows both binary and multiple values. A function $f\!\in\!\mathcal{F}:\mathcal{X\!\mapsto\!Y}$ represents a hypothesis in a space of hypotheses $\mathcal{F}$, of which the prediction for one instance $\bm{x}$ is denoted by $f(\bm{x})$ or $\hat{y}$.
@@ 3.1 Model fairness assessment from a manifold perspective
Given the dataset $S\!=\!\{(\mathop{\breve{\bm{x}}}_{i},\mathop{\bm{a}}_{i},y_{i})|i\in[n]\}$ composed of instances including sensitive attributes, here we denote one instance by $\bm{x}=(\mathop{\breve{\bm{x}}},\mathop{\bm{a}})=[x_{1},...,x_{n_{x}},a_{1},...,a_{n_{a}}]^{\mathsf{T}}$ for clarity, where $n_{a}$ is the number of sensitive/protected attributes and $n_{x}$ is that of unprotected attributes in $\bm{x}$. In this paper, we introduce new fairness measures in scenarios for multiple sensitive attributes with multiple possible values. Note that the proposed fairness measure in this paper is extended from our previous work—a fairness measure in scenarios for sensitive attributes with binary values [34].
3.1.1 Distance between sets for sensitive attributes with binary values, from our previous work [34]
Inspired by the principle of individual fairness—similar treatment for similar individuals, if we view the instances (with the same sensitive attributes) as data points on certain manifolds, the manifold representing members from the marginalised/unprivileged group(s) is supposed to be as close as possible to that representing members from the privileged group. To measure the fairness with respect to the sensitive attribute, we have proposed a fairness measure that is inspired by ‘the distance of sets’ introduced in mathematics. For a certain bi-valued sensitive attribute $a_{i}\!\in\!\mathcal{A}_{i}\!=\!\{0,1\}$, $S$ can be divided into two subsets $S_{1}\!=\!\{(\bm{x},y)\!\in S|a_{i}\!=\!1\}$ and $\bar{S}_{1}\!=\!S\setminus S_{1}\!=\!\{(\bm{x},y)\!\in S|a_{i}\!\neq\!1\}$, where $a_{i}\!=\!1$ means the corresponding instance is a member from the privileged group. Then given a specific distance metric $\mathbf{d}(\cdot,\cdot)$ [^1] Here we use the standard Euclidean metric. In fact, any two metrics $\mathbf{d}_{1},\mathbf{d}_{2}$ derived from norms on the Euclidean space $\mathbb{R}^{d}$ are equivalent in the sense that there are positive constants $c_{1},c_{2}$ such that $c_{1}\mathbf{d}_{1}(x,y)\leqslant\mathbf{d}_{2}(x,y)\leqslant c_{2}\mathbf{d}_{1}(x,y)$ for all $x,y\in\mathbb{R}^{d}$. on the feature space, our previous distance between these two subsets (that is, $S_{1}$ and $\bar{S}_{1}$) is defined by

### v2

@@ Abstract
Discrimination mitigation within machine learning (ML) models could be complicated because multiple factors may be interwoven hierarchically and historically. Yet few existing fairness measures can capture the discrimination level within ML models in the face of multiple sensitive attributes (SAs). To bridge this gap, we propose a fairness measure based on distances between sets from a manifold perspective, named as ‘Harmonic Fairness measure via Manifolds (HFM)’ with two optional versions, which can deal with a fine-grained discrimination evaluation for several SAs of multiple values. Because directly computing HFM may be costly, to accelerate its subprocedure—the computation of distances of sets, we further propose two approximation algorithms named ‘Approximation of distance between sets for one sensitive attribute with multiple values (ApproxDist)’ and ‘Approximation of extended distance between sets for several sensitive attributes with multiple values (ExtendDist)’ to respectively resolve bias evaluation of one single SA with multiple values and that of several SAs with multiple values. Moreover, we provide an algorithmic effectiveness analysis for ApproxDist under certain assumptions to explain how well it could work. The empirical results demonstrate that our proposed fairness measure HFM is valid and approximation algorithms (i.e. ApproxDist and ExtendDist) are effective and efficient.
@@ 1 Introduction
As techniques of machine learning (ML) and deep learning (DL) are flourishingly developed and ML/DL systems are widely deployed in real life nowadays, concerns about the underlying discrimination hidden in these models has grown, particularly in high-stakes domains such as healthcare, recruitment, and jurisdiction [1], where equity for all stakeholders is pivotal to prevent unjust outcomes, akin to a discriminatory Matthew Effect. It is important to prevent ML models from perpetuating or exacerbating inappropriate human prejudices not only for model performance but also for societal welfare. Effectively addressing and eliminating discrimination usually requires a comprehensive grasp of its occurrence, causes, and mechanisms. For instance, a case involving a person changing their gender for lower car insurance rates highlights the complexity of fairness in ML.
@@
Although the impressive practical advancements of ML and DL thrive on abundant data, their trustworthiness and equity heavily hinge on data quality. In fact, one of the primary sources of unfairness identified in the existing literature is biases from the data, possibly collected from various sources such as device measurements and historically biassed human decisions [2]. Moreover, the challenge of data imbalance often looms in human-sensitive domains, amplifying concerns of discrimination and bias propagation in ML models. As a result, misformed model training would amplify imbalances and biases in data, with wide-reaching societal implications. For example, optimising aggregated prediction errors can advantage privileged groups over marginalised ones. In addition, missing data such as instances or values may introduce disparities between the dataset and the target population, leading to biassed results as well. Therefore, in order to ensure fairness and mitigate biases, it is crucial to properly address data imbalance and prevent ML models from perpetuating or even exacerbating inappropriate human prejudices.
@@
To mitigate bias within ML models, the very first step is to promptly recognise its occurrence. However, promptly detecting discrimination fully, truly, and faithfully is not quite easy because of plenty of factors interwoven with each other. First, learning algorithms might yield unfair outcomes even with purely clean data due to proxy attributes for sensitive features or biassed algorithmic objectives. For instance, the educational background of one person might be a proxy attribute for those born in families with a preference for boys. Second, the existence of multiple sensitive attributes (SAs) and their interaction with each other highlights the complexity of bias tackling, like one member from a marginalised group could become one of the majority concerning another factor, or vice versa. Third, dynamic changes and historical factors may need to be taken into account, as bias hidden in data, data imbalance, and present decisions may interweave, causing interrelated impacts and vicious circles. Despite many fairness measures that have been proposed to facilitate bias mitigation, most of them mainly focus on one single SA or ones with binary values, and few could handle bias appropriately when facing multiple SAs with even multiple values. Therefore, it motivates us to investigate a proper tool to deal with bias in such aforementioned scenarios.
@@
In this paper, we investigate the possibility of assessing the discrimination level of ML models in the presence of several SAs with multiple values. To this end, we introduce a novel fairness measure from a manifold perspective, named ‘Harmonic Fairness measure via Manifolds (HFM)’, with two optional versions (that is, maximum HFM and average HFM). However, the direct calculation of HFM lies on a core distance between two sets, which might be pretty costly. Therefore, we further propose two approximation algorithms that quickly estimate the distance between sets, named as ‘Approximation of distance between sets for one sensitive attribute with multiple values (ApproxDist)’ and ‘Approximation of extended distance between sets for several sensitive attributes with multiple values (ExtendDist)’ respectively, in order to speed up the calculation, accelerate the bias evaluation, and broaden its practical applicability. Furthermore, we also investigate their algorithmic properties under certain reasonable assumptions, in other words, how effective they could be in achieving the approximation goal. Our contribution in this work is four-fold:
@@
  • We propose a fairness measure named HFM that could reflect the discrimination level of classifiers even simultaneously facing several SAs with multiple values. Note that HFM has two optional versions, of which both are built upon a concept of distances between sets from the manifold perspective.
@@
  • We propose two approximation algorithms (that is, ApproxDist and ExtendDist) that accelerate the estimation of distances between sets, to mitigate the disadvantage of costly direct calculation of HFM.
@@
  • We further investigate the algorithmic effectiveness of ApproxDist under certain assumptions and provide detailed explanations.
@@ 2.1 Techniques to enhance fairness
Existing mechanisms to mitigate biases and enhance fairness in ML models can typically be divided into three types: pre-processing, in-processing, and post-processing mechanisms, based on when manipulations are applied during model training pipelines. Particularly, recent work on in-processing fairness for DL models mainly falls under two types of approaches: constraint-based and adversarial learning methods [3]. Constraint-based methods usually incorporate fairness metrics directly into the model optimisation objectives as constraints or regularisation terms. For instance, Zemel et al. [4], the pioneer in this direction, put demographic parity constraints on model predictions. Subsequent work also includes using approximations [5] or modified training schemes [6] to improve scalability. Adversarial methods intend to learn representations as fairly as possible by removing sensitive attribute information. In such procedures, additional prediction heads may be introduced for attribute subgroup predictions, and the information concerning sensitive attributes would be removed through inverse gradient updating [7, 8] or disentangling features [9, 10, 11, 12]. Other fairness enhancing techniques include data augmentations [13], sampling [14, 15], data noising [16], dataset balancing with generative methods [17, 18, 19], and reweighting mechanisms [20, 21]. Recently, mixup operations [22, 23, 3] are adopted to enhance fairness by blending inputs across subgroups [24, 25]. Yet most of these studies focus on protecting one sensitive attribute and are hardly able to handle several sensitive attributes all at once. And multi-attribute fairness protection remains relatively rarely explored.
@@ 2.2 Existing fairness metrics and multi-attribute fairness protection
The well-known fairness metrics are generally divided into group fairness—such as demographic parity (DP), equality of opportunity (EO), and predictive quality parity (PQP)—and individual fairness [26, 27, 28, 29, 30]. The former mainly focuses on statistical/demographic equality among groups defined by sensitive attributes, while the latter cares more about the principle that ‘similar individuals should be evaluated or treated similarly.’ However, satisfying fairness metrics all at once is hard to achieve because they are usually not compatible with each other [31]. In practice, it may need to deliberate on the choice of the specified distance in individual fairness [29, 26]. Moreover, the three commonly used group fairness measures (that is, DP, EO, and PQP) can only deal with one single sensitive attribute with binary values. Although extending them to scenarios of one sensitive attribute with multiple values is possible like statistical parity [32, 33], they are still limited when facing several sensitive attributes at the same time. Recent work includes a newly proposed fairness measure named discriminative risk (DR) [34] that is capable of capturing bias from both individual and group fairness aspects and two fairness frameworks (that is, InfoFair [35] and MultiFair [3]) to deliver fair predictions in face of multiple sensitive attributes. Yet these two fairness frameworks are not quantitative measures that directly evaluate the discrimination level of ML models.
@@ 3 Methodology
In this section, we formally study the measurement of fairness from a manifold perspective. Some standard notations that we use in this paper is listed in Table I.
TABLE I: Mathematical notation summary
Notation Meaning
$x$ Scalar (italic lowercase letter)
$\bm{x}$ Vector (bold lowercase letter)
$X$ Matrix or set (italic uppercase letter)
$\mathsf{X}$ Random variable (serif uppercase letter)
$\mathbb{R}$ Real numbers
$\mathbb{Z}$, $\mathbb{Z}^{+}$ Integers, and positive integers
$\mathbb{P}(\cdot)$ Probability measure
$\mathbb{E}(\cdot)$ Expectation
$\mathbb{V}(\cdot)$ Variance of a random variable
$\mathbb{I}(\cdot)$ Indicator function
$\mathcal{F}$ Hypothesis space
$f(\cdot)$ Models in one hypothesis space
$[n]$ The set $\{1,2,...,n\}$ for brevity
Furthermore, we define:
  • $S=\{(\bm{x}_{i},y_{i})\}_{i=1}^{n}$ denotes a dataset with i.i.d. instances drawn from a feature-label space $\mathcal{X}\times\mathcal{Y}$ based on an unknown distribution.
  • The feature (or input) space $\mathcal{X}$ is arbitrary, while the label (or output) space $\mathcal{Y}\!=\!\{1,2,\ldots,n_{c}\}$ ($n_{c}\geqslant 2$) is finite, supporting both binary and multi-class classification.
  • For dataset $S$ containing sensitive attributes (SAs), each instance is represented as $\bm{x}\!\triangleq\!(\mathop{\breve{\bm{x}}},\mathop{\bm{a}})$, where:
    • $\mathop{\breve{\bm{x}}}=[x_{1},x_{2},\ldots,x_{n_{x}}]^{\mathsf{T}}$ represents non-sensitive features, and $n_{x}$ is the number of non-sensitive features;
    • $\mathop{\bm{a}}=[a_{1},a_{2},\ldots,a_{n_{a}}]^{\mathsf{T}}$ represents SAs, and the number of SAs $n_{a}\geqslant 1$, allowing multiple attributes; and
    • $a_{i}\in\mathbb{Z}_{+}$ for the $i$-th SA ($1\leqslant i\leqslant n_{a}$), allowing both binary and multiple values.
  • A hypothesis function $f\in\mathcal{F}:\mathcal{X}\mapsto\mathcal{Y}$ maps from the feature space to the label space, where:
    • $\mathcal{F}$ is the hypothesis space, and
    • $f(\bm{x})$ or $\hat{y}$ denotes the prediction for instance $\bm{x}$.
@@ 3.1 Model fairness assessment from a manifold perspective
Given the dataset $S\!=\!\{(\mathop{\breve{\bm{x}}}_{i},\mathop{\bm{a}}_{i},y_{i})\mid i\!\in\![n]\}$ composed of instances including SAs, here we denote one instance by $\bm{x}\!=\!(\mathop{\breve{\bm{x}}},\mathop{\bm{a}})\!=\![x_{1},...,x_{n_{x}},a_{1},...,a_{n_{a}}]^{\mathsf{T}}$ for clarity, where $n_{a}$ is the number of sensitive/protected attributes and $n_{x}$ is that of non-sensitive/unprotected attributes in $\bm{x}$. In this paper, we introduce new fairness measures in scenarios for several SAs with multiple possible values. Note that the proposed fairness measure here is extended from our previous work—a fairness measure in scenarios for SAs with binary values [36].
3.1.1 Distance between sets for one bi-valued SA, from our previous work [36]
Inspired by the principle of individual fairness—similar treatment for similar individuals, if we view the instances (with the same SAs) as data points on certain manifolds, the manifold representing members from the marginalised/unprivileged group(s) is supposed to be as close as possible to that representing members from the privileged group. To measure the fairness with respect to the SA, we have proposed a fairness measure that is inspired by ‘the distance of sets’ introduced in mathematics.111 The distance between sets used here is known as the Hausdorff distance (HD), denoted as $\mathbf{d}_{H}$. It measures the distance between two non-empty subsets (namely $X$ and $Y$) in a metric space $(\mathcal{M},\mathbf{d})$, it is defined as $\footnotesize\mathbf{d}_{H}\triangleq\max\big{\{}\sup_{x\in X}\inf_{y\in Y}\mathbf{d}(x,y),\\ \sup_{y\in Y}\inf_{x\in X}\mathbf{d}(x,y)\big{\}}\,.$ When dealing with discrete sets, we may replace supremum and infimum with maximum and minimum respectively. We chose it over any other distance for two reasons. One is that if a whole data space is divided by one bi-valued SA, the obtained subspaces are assumed to be disjoint, and the whole data space is viewed as a Hausdorff space, a premise when using the HD. The other is that the HD is the greatest of all the distances from one point in one set to the closest point in the other set, which helps us not be misled by underestimated or overly optimistic results, considering that we were trying to capture the discrimination within. For a certain bi-valued SA $a_{i}\!\in\!\mathcal{A}_{i}\!=\!\{0,1\}$, $S$ can be divided into two subsets $S_{1}\!=\!\{(\bm{x},y)\!\in S\mid a_{i}\!=\!1\}$ and $\bar{S}_{1}\!=\!S\setminus S_{1}\!=\!\{(\bm{x},y)\!\in S\mid a_{i}\!\neq\!1\}$, where $a_{i}\!=\!1$ means the corresponding instance is a member of the privileged group. Then given a specific distance metric $\mathbf{d}(\cdot,\cdot)$ [^2] Here we use the standard Euclidean metric. In fact, any two metrics $\mathbf{d}_{1},\mathbf{d}_{2}$ derived from norms on the Euclidean space $\mathbb{R}^{d}$ are equivalent in the sense that there are positive constants $c_{1},c_{2}$ such that $c_{1}\mathbf{d}_{1}(x,y)\leqslant\mathbf{d}_{2}(x,y)\leqslant c_{2}\mathbf{d}_{1}(x,y)$ for all $x,y\in\mathbb{R}^{d}$. on the feature space, our previous distance between these two subsets (that is, $S_{1}$ and $\bar{S}_{1}$) is defined by